Question- The displacement of a particle at any instant is given by x= 8t2 – 3t3.  Calculate the average velocity in the interval from t=0 to t=2s and in the time interval from t=0 to t=3s,if x is measured in m. A boy runs along a straight path for the first half distance with a velocity v and the second half distance with velocity v . Find the average velocity of the boy. 

Grade:11

1 Answers

Eshan
5 years ago
Dear student,

Please ask only one question per post. Answer to first question is given below-

x=8t^2-3t^3
Hence the displacement from t=0 to t=2s isx(2)-x(0)=8(2)^2-3(2)^3=8m
Therefore, the average velocity in the period=\dfrac{8m}{2s}=4m/s